Flounder Distribution markets CDs of the performing artist Fishe. At the beginning of October, Flounder had in beginning inventory 2,800 of Fishe’s CDs with a unit cost of $6. During October, Flounder made the following purchases of Fishe’s CDs. Oct. 3 3,500 @ $7 Oct. 19 4,200 @ $9 Oct. 9 4,900 @ $8 Oct. 25 5,600 @ $10 During October, 15,260 units were sold. Flounder uses a periodic inventory system.

Calculate weighted-average cost per unit. (Round answer to 2 decimal places, e.g. 2.25.) Weighted-average cost per unit
Answer #1

2800 * 6 = 16,800

3500 * 7 = 24,500

4200 * 9 = 37,800

4900 * 8 = 39,200

5600 * 10 = 56,000

Total Cost = 174,300 / 21,000

Weighted average cost per unit = 8.30